OpenStudy (anonymous):

how do I factor 2x^2+3x-5

OpenStudy (the_fizicx99):

Again, find a product that when you multiply 2 * -5 gives you -10 and when added gives you +3 5 * -2 = -10 5 + -2 = 3 2x^2 + 5x - 2x - 5 x(2x + 5) -1(2x + 5) (2x + 5)(x - 1)

OpenStudy (anonymous):

okay i did all of that what do i do next

jimthompson5910 (jim_thompson5910):

(2x + 5)(x - 1) is the fully factored form, so once you get there, you're done

OpenStudy (anonymous):

okay thank you i have another one

jimthompson5910 (jim_thompson5910):

ok go ahead

OpenStudy (anonymous):

okay 6x^2+11x+3 factor

jimthompson5910 (jim_thompson5910):

that is in the form ax^2 + bx + c where a = 6 b = 11 c = 3 see how I'm getting this?

OpenStudy (anonymous):

yues and no the x ?? confuse me

jimthompson5910 (jim_thompson5910):

well we don't know what x is in 6x^2+11x+3, so we just leave it as x for now

jimthompson5910 (jim_thompson5910):

same for ax^2 + bx + c

OpenStudy (anonymous):

okay so what do i do next

jimthompson5910 (jim_thompson5910):

what is the value of a*c ?

OpenStudy (anonymous):

a: 6 and c: 3 right

jimthompson5910 (jim_thompson5910):

good, so a*c = ???

OpenStudy (anonymous):

6^3 right ??

jimthompson5910 (jim_thompson5910):

a*c means a times c

jimthompson5910 (jim_thompson5910):

a is 6, c is 3, so a*c = 6*3 = ??

OpenStudy (anonymous):

yeah that is what i put okay you can go on to the next step now

jimthompson5910 (jim_thompson5910):

what is 6*3

OpenStudy (anonymous):

18 right

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so you need to find two numbers that multiply to 18 AND add to 11 (11 is the middle coefficient)

OpenStudy (anonymous):

random numbers

jimthompson5910 (jim_thompson5910):

two whole numbers

jimthompson5910 (jim_thompson5910):

what two whole numbers both multiply to 18 AND add to 11

OpenStudy (anonymous):

i don't really know

jimthompson5910 (jim_thompson5910):

what are some ways to multiply to 18?

jimthompson5910 (jim_thompson5910):

using whole numbers

OpenStudy (anonymous):

1,2,3,8 right

jimthompson5910 (jim_thompson5910):

3*6 is one way to multiply to 18, what's another?

OpenStudy (anonymous):

2*9 right

jimthompson5910 (jim_thompson5910):

good, what happens when you add 2 and 9?

OpenStudy (anonymous):

you get 11 ohhh i see now okay move on

jimthompson5910 (jim_thompson5910):

so the numbers 2,9 both multiply to 18 and add to 11

jimthompson5910 (jim_thompson5910):

that means we break up 11x into 2x + 9x

jimthompson5910 (jim_thompson5910):

6x^2+11x+3 turns into 6x^2+2x+9x+3

OpenStudy (anonymous):

okay what do we do from there

jimthompson5910 (jim_thompson5910):

you factor by grouping

jimthompson5910 (jim_thompson5910):

have you done that before?

OpenStudy (anonymous):

no i don't think so

jimthompson5910 (jim_thompson5910):

first you group the terms in pairs like this 6x^2+2x+9x+3 (6x^2+2x)+(9x+3)

jimthompson5910 (jim_thompson5910):

What can you factor from (6x^2+2x)

OpenStudy (anonymous):

6x^2

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

what is the GCF of 6x^2 and 2x

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

6 is not a factor of 2

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

is a factor of both 2 and 6, good

jimthompson5910 (jim_thompson5910):

what else is a factor of 6x^2 and 2x

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

what do they share together?

jimthompson5910 (jim_thompson5910):

other than a factor of 2

OpenStudy (anonymous):

6, and 12

jimthompson5910 (jim_thompson5910):

x^2 is the same as x*x, agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so let's find the prime factorization of 6x^2 and 2x